OpenStudy (anonymous):

Use the slope and y-intercept, if possible, to graph the line whose equation is given. y=2x-1

OpenStudy (anonymous):

well thats easy. you use y= mx + b your 2x is your m and your b is -1

OpenStudy (anonymous):

Meaining? I am confused on this.

OpenStudy (anonymous):

the eq. y= 2x -1 is in the slope-intercept form (y=mx +b) the coeff of x is the slope or the rate at which a line is going up or coming down, reading the graph from right to left. so the slope of this equation/line is 2 or more precisely 2 units of y for every one (1) unit of x. you usually hear it as y/x or rise/run. the -1 is the y intercept that is the point at which the equation's line will cross the y axis. so the line is going to go throughh -1 on the y axis, so now find some data points for this eq,. make table x y let x=0 0 what is y? 1 " plug it into the eq. y=2x-1............................y=-1 so you have one point now let x=1 what is y now?...................y=1 now you another point.......and you can draw thr line of the eq. the line ought to pass thru y, -1 if it doesn't something is wrong.